OpenStudy (anonymous):

x-10y=-3 & 7y-4x=-21. Please show me to find an ordered pair.

OpenStudy (amistre64):

you want a solution to both equations right?

OpenStudy (amistre64):

in other words, these are two lines the cross at one point... and find that point right?

OpenStudy (amistre64):

ill show you the elimination method. we start by stacking them on top of each other and we are going to "eliminate" one of the variables.... its cruel I know, but its gotta be done :) -4x +7y =-21 x -10y =-3 <- mulitiply this one by (4) -4x +7y =-21 4x -40y =-12 ; now add the columns together -------------- 0 -33y = -33 ^ see what happened here? we eliminated the "x"s now we can solve this for "y"

OpenStudy (amistre64):

-33y = -33 ; divide both sides by -33 to get: -33y/-33 = -33/-33 ; anything over itself = 1 y = 1 Since we know a "solid" value for "y"; plug it into any of the equations and solve for the "x"

OpenStudy (amistre64):

x -10y =-3 x -10(1) = -3 x - 10 = -3 ; add 10 to both sides x -10+10 = -3+10 x-0 = 7 x = 7 your ordered pair is (7,1)

OpenStudy (anonymous):

I wll have to look at this. This is something they do not show us in our books. Thank you very much.

OpenStudy (amistre64):

youre welcome...

OpenStudy (anonymous):

I have some more problems like thisand will use this to see how it works out. Again Thank you.